OpenStudy (anonymous):

If Γ |= φ and Δ |= φ, then it must be the case that Γ ∩ Δ |= φ. True or False?

OpenStudy (anonymous):

false unless Γ ∩ Δ l = 0

OpenStudy (anonymous):

Can you explain?

OpenStudy (anonymous):

I was thinking along the lines of 1+2=3, and 1+2=3, then 1+2+1+2=3 doesn't work but 0+0+0+0=0...not sure if I'm thinking of it right though..

OpenStudy (anonymous):

ok